OpenStudy (anonymous):

sqrt(30) + 6sqrt(10) = ? (has to be in standard form example 2sqrt(7) )

OpenStudy (anonymous):

calculator 24.450891536061936134569697828008

OpenStudy (anonymous):

thanks sorry it needs to be expressed in standard form.

OpenStudy (anonymous):

what u mean standard form?

OpenStudy (anonymous):

I revised the question

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \sqrt{30} = \sqrt{3*10}\] \[\Large \sqrt{30} = \sqrt{3}*\sqrt{10}\]